OpenStudy (anonymous):

36^(3c)=sqrt6^(8c+6)

jimthompson5910 (jim_thompson5910):

\[\Large 36^{3c} = \sqrt{6^{8c+6}}\] \[\Large 36^{3c} = \left(6^{8c+6}\right)^{\frac{1}{2}}\] \[\Large (6^2)^{3c} = \left(6^{8c+6}\right)^{\frac{1}{2}}\] \[\Large 6^{2*3c} = 6^{\frac{1}{2}(8c+6)}\] \[\Large 6^{6c} = 6^{\frac{1}{2}(8c)+\frac{1}{2}(6)}\] \[\Large 6^{6c} = 6^{4c+3}\] \[\Large 6c = 4c+3\] \[\Large 6c-4c = 3\] \[\Large 2c = 3\] \[\Large c = \frac{3}{2}\]